IP angles scan experiment - 05 March 2007

HER Single beam: Push current

13:50 XP and XP at optimum values from scans
13:52 Start trickling up
14:12 I_HER = 785 mA. ST BPMs and HIPP BPMs are NOT yet saturated. Stopping here for Thermal limit.
14:15 YP still at 1.1 mrad. However, XP down by 700 urad compared to earlier value. Re-adjusted XP.
14:22 HER dumped, for LER scan